an apartment building casts a shadow. From the tip of the shadow to the top of the building is 100 m. The tip of the shadow is 72 m from the base of the building? How tall is the building?
thnakyou

Answer by stanbon(48558)

an apartment building casts a shadow. From the tip of the shadow to the top of the building is 100 m. The tip of the shadow is 72 m from the base of the building? How tall is the building?
--------------------
Draw the picture and label the height of the building "x".
-----------------------
Use Pythagoras or use trigonometry if you know it.
100^2 = 72^2 + x^2
x^2 = 10000 - 5184
x^2 = 4816
x = 69.40 meters
=======================
Cheers,
Stan H.